# Simple homotopy types of even dimensional manifolds

### v2

@@ 3.2. Realising via the surgery exact sequence
First assume that $n=2q$ is even. Since $f$ is $q$-connected, $\sigma_{s}(\eta_{h}([N,f]))$ is defined as a form on $\ker(H_{q}(f))\cong H_{q+1}(\operatorname{Cyl}_{f},N)$, which is trivial (because $f$ is a homotopy equivalence), and it is equipped with a certain stable basis. The stable basis of $H_{q+1}(\operatorname{Cyl}_{f},N)=0$ is chosen such that with this choice $\tau(\operatorname{Cyl}_{f},N)=0$ (where $H_{i}(\operatorname{Cyl}_{f},N)=0$ is equipped with the canonical stable basis if $i\neq q+1$). If $H_{q+1}(\operatorname{Cyl}_{f},N)$ were also equipped with the canonical stable basis, then the Whitehead torsion of $(\operatorname{Cyl}_{f},N)$ would be equal to $\tau(f)$. Therefore the transition matrix between the chosen and the standard stable basis (when both are regarded as bases of $\mathbb{Z}G^{k}$ for some $k\geq 0$) has Whitehead torsion $(-1)^{q}\tau(f)$. In $L^{s}_{n}(\mathbb{Z}G,w)$ the same element $\sigma_{s}(\eta_{h}([N,f]))$ is also represented by a standard hyperbolic form with a basis such that the transition matrix between the chosen and the standard basis has Whitehead torsion $(-1)^{q}\tau(f)$.
@@
The image of $\pi(\tau(f))$ in $L^{s}_{n}(\mathbb{Z}G,w)$ is represented by a standard hyperbolic form with a basis with the property that if $x$ denotes the Whitehead torsion of the transition matrix between the chosen and the standard basis, then $x\in\mathcal{J}_{n}(G,w)$ and $\pi(x)=\pi(\tau(f))$ (see [Shaneson-GxZ, p. 312]). In particular, the representative of $\sigma_{s}(\eta_{h}([N,f]))$ constructed above also represents $\varrho(\widehat{\tau}([N,f]))$, noting that $\pi((-1)^{q}\tau(f))=\pi(\tau(f))$, because $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is $2$-torsion. Hence the square commutes if $n$ is even.
@@
Next consider the case when $n=2q+1$ is odd. Let $U$ denote a tubular neighbourhood of a disjoint union of embeddings $S^{q}\rightarrow N$ representing a generating set of $\ker(\pi_{q}(f))$ and let $N_{0}=N\setminus\operatorname{Int}U$. We identify $H_{q}(\partial U)$ with the standard hyperbolic form, then $H_{q+1}(U,\partial U)$ (more precisely, its image under the boundary map) corresponds to the standard lagrangian. We can assume that $f\big{|}_{U}$ is constant, so $f\big{|}_{N_{0}}$ is a map of pairs $(N_{0},\partial U)\rightarrow(M,\ast)$. Then $\ker(H_{q+1}(N_{0},\partial U)\rightarrow H_{q+1}(M,\ast))\cong H_{q+2}(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}})$ determines another lagrangian in $H_{q}(\partial U)$. We equip $H_{q}(\partial U)$ and $H_{q+1}(U,\partial U)$ with their standard bases, and $H_{q+2}(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}})$ with a stable basis such that $\tau(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}})=0$. Then $\sigma_{s}(\eta_{h}([N,f]))$ is represented by the formation $(H_{q}(\partial U);H_{q+1}(U,\partial U),H_{q+2}(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}}))$. Since $f$ is a homotopy equivalence, we can take the empty generating set for $\ker(\pi_{q}(f))$. Then $H_{q}(\partial U)=0$, so we have the trivial formation, with the standard basis on the ambient form and on the first lagrangian. The stable basis on the second lagrangian, $H_{q+2}(\operatorname{Cyl}_{f},N)$, is chosen such that $\tau(\operatorname{Cyl}_{f},N)=0$. Therefore the transition matrix between this stable basis and the standard one has Whitehead torsion $(-1)^{q+1}\tau(f)$. In $L^{s}_{n}(\mathbb{Z}G,w)$ the same element $\sigma_{s}(\eta_{h}([N,f]))$ is also represented by a formation on the standard hyperbolic form given by the standard lagrangians, such that the ambient form and the first lagrangian are equipped with their standard bases, and for the second lagrangian the transition matrix between the chosen and the standard basis has Whitehead torsion $(-1)^{q+1}\tau(f)$.
@@
The image of $\pi(\tau(f))$ in $L^{s}_{n}(\mathbb{Z}G,w)$ is represented by a formation on the standard hyperbolic form given by the standard lagrangians, such that the ambient form and the first lagrangian are equipped with their standard bases, and if $x$ denotes the Whitehead torsion of the transition matrix between the chosen and the standard basis of the second lagrangian, then $x\in\mathcal{J}_{n}(G,w)$ and $\pi(x)=\pi(\tau(f))$. Again we use that we can ignore signs because $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is $2$-torsion. So the representative of $\sigma_{s}(\eta_{h}([N,f]))$ constructed above also represents $\varrho(\widehat{\tau}([N,f]))$, showing that the square also commutes if $n$ is odd. ∎
@@
  1. (a)
    $\operatorname{Im}\widehat{\tau}=\varrho^{-1}(\operatorname{Im}\sigma_{s})$.
  2. (b)
    $\operatorname{Im}\psi\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{s})$.
@@
First we prove (a). The commutativity of the square in Proposition 3.7 implies that $\operatorname{Im}\widehat{\tau}\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{s})$. For the other direction assume that $x\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ and $\varrho(x)=\sigma_{s}(y)$ for some $y\in\mathcal{N}(M)$. Then $\sigma_{h}(y)=F\circ\sigma_{s}(y)=F\circ\varrho(x)=0$, so $y=\eta_{h}(v)$ for some $v\in\mathcal{S}^{h}(M)$. Then $\varrho(x-\widehat{\tau}(v))=\varrho(x)-\sigma_{s}(\eta_{h}(v))=\varrho(x)-\sigma_{s}(y)=0$, so $x-\widehat{\tau}(v)=\psi(z)$ for some $z\in L^{h}_{n+1}(\mathbb{Z}G,w)$. So by Proposition 3.6 $x=\widehat{\tau}(v)^{z}=\widehat{\tau}(v^{z})\in\operatorname{Im}\widehat{\tau}$.
For  (b), since $\operatorname{Id}_{M}$ has vanishing surgery obstruction, we have $\operatorname{Im}\psi=\ker\varrho\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{s})$. ∎
@@
Let $x\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$. If $x\in\operatorname{Im}\psi$, then there is a $\operatorname{CAT}$ $n$-manifold $N$ and a homotopy equivalence $f\colon N\to M$ such that $\pi(\tau(f))=x$.
@@
By Corollary 3.9, $x=\widehat{\tau}([N,f])$ for some $[N,f]\in\mathcal{S}^{h}(M)$. ∎
@@
There is a common generalisation of the two realisation techniques used in Section 3. Let $L_{n}^{s,\tau}(\mathbb{Z}G,w)$ denote the surgery obstruction groups defined in [Kreck-preprint, Section 4]. Similarly to $L_{n}^{s}(\mathbb{Z}G,w)$, the elements of $L_{n}^{s,\tau}(\mathbb{Z}G,w)$ are represented by based nonsingular forms/formations, but, unlike the elements of $L_{n}^{s}(\mathbb{Z}G,w)$, they are not assumed to be simple. Thus there is a natural forgetful map $L_{n}^{s,\tau}(\mathbb{Z}G,w)\to L_{n}^{h}(\mathbb{Z}G,w)$ and $L_{n}^{s}(\mathbb{Z}G,w)$ is the kernel of a map $L_{n}^{s,\tau}(\mathbb{Z}G,w)\to\operatorname{Wh}(G,w)$ (see [Kreck-preprint]). There is also an action of $\operatorname{Wh}(G,w)$ on $L_{n}^{s,\tau}(\mathbb{Z}G,w)$ (which determines a map $\operatorname{Wh}(G,w)\to L_{n}^{s,\tau}(\mathbb{Z}G,w)$ by acting on 00) given by changing the basis; this action is transitive on the fibers of the $L_{n}^{s,\tau}(\mathbb{Z}G,w)\to L_{n}^{h}(\mathbb{Z}G,w)$.
@@
Wall’s construction for realising elements of $L_{n}^{s}(\mathbb{Z}G,w)$  [Wall-SCM] can also be applied to $L_{n}^{s,\tau}$, but it will only produce homotopy equivalences, not simple homotopy equivalences. The Whitehead torsion of the resulting homotopy equivalence is given by the previously mentioned map $L_{n}^{s,\tau}(\mathbb{Z}G,w)\to\operatorname{Wh}(G,w)$, and in fact the image of this map is in $\mathcal{J}_{n}(G,w)$.
@@
The realisation of $L_{n}^{h}(\mathbb{Z}G,w)$, and the map $W_{h}$, can be regarded as a special case of this construction, by first choosing a basis for the form/formation representing an element of $L_{n}^{h}(\mathbb{Z}G,w)$, equivalently, choosing a lift in $L_{n}^{s,\tau}(\mathbb{Z}G,w)$. As the choice of basis is not unique, the homotopy equivalence we obtain is not unique, and its Whitehead torsion (the value of $\widehat{\tau}$) is only well-defined in the quotient $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$. The realisation part of the $s$-cobordism theorem is also a special case via the map $\operatorname{Wh}(G,w)\to L_{n}^{s,\tau}(\mathbb{Z}G,w)$. The Whitehead torsion of the homotopy equivalence induced by the $h$-cobordism we get is of course in $\mathcal{I}_{n}(G,w)$.
In the rest of this section we will consider the orientable ($w\equiv 1$) case, and omit $w$ from the notation.
Proposition 3.12.
Let $G_{m}\coloneqq C_{\infty}\times C_{m}$, with $m\geq 2$. For every even integer $n=2k\geq 0$, the map
$\psi\colon L_{n+1}^{h}(\mathbb{Z}[G_{m}])\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G_{m}))$
is surjective.
Proof.
We verify that the forgetful map $F\colon L^{s}_{2k}(\mathbb{Z}G_{m})\to L^{h}_{2k}(\mathbb{Z}G_{m})$ is injective. The conclusion then follows from the exact sequence (3.1). First we apply Shaneson splitting to the domain and the codomain, to obtain a commutative diagram whose rows are split short exact sequences and whose vertical maps are the forgetful maps [Ra86].
00{0}${L^{s}_{2k}(\mathbb{Z}C_{m})}$${L^{s}_{2k}(\mathbb{Z}G_{m})}$${L^{h}_{2k-1}(\mathbb{Z}C_{m})}$00{0}00{0}${L^{h}_{2k}(\mathbb{Z}C_{m})}$${L^{h}_{2k}(\mathbb{Z}G_{m})}$${L^{p}_{2k-1}(\mathbb{Z}C_{m})}$${0.}$$\scriptstyle{F}$
By the five lemma it suffices to show that the left and right vertical maps are injective. In [Bak-computation-L-groups], Bak gave computations of $L_{i}^{X}(\mathbb{Z}G)$ for $X\in\{s,h,p\}$ and $G$ a finite group whose 2-hyperelementary subgroups are abelian, which certainly holds for finite cyclic groups. See also [Bak-odd-L-groups-of-odd-order-groups, Bak-even-L-groups-of-odd-order-groups] First we note that, by [Bak-computation-L-groups, Theorem 8], the forgetful map $L^{h}_{2k-1}(\mathbb{Z}C_{m})\to L^{p}_{2k-1}(\mathbb{Z}C_{m})$ is injective. So the right vertical map is injective.
To prove that the left vertical map is injective we also use [Bak-computation-L-groups]. For a finite group $G$ whose 2-Sylow subgroup $G_{2}$ is normal and abelian, Bak defined $r_{2}:=\operatorname{rk}H^{1}(C_{2};\operatorname{Wh}(G))$. On [Bak-computation-L-groups, p. 386], he noted that if $G_{2}$ is cyclic, as in our case $G=C_{m}$, then $r_{2}=0$. Therefore by the sequence (3.1), as shown on [Bak-computation-L-groups, p. 390], it follows that $L_{2k}^{s}(\mathbb{Z}C_{m})\to L_{2k}^{h}(\mathbb{Z}C_{m})$ is injective. Hence $F\colon L_{2k}^{s}(\mathbb{Z}G_{m})\to L_{2k}^{h}(\mathbb{Z}G_{m})$ is injective as desired. ∎

### v3

@@ 3.2. Realising via the surgery exact sequence
First assume that $n=2q$ is even. Since $f$ is $q$-connected, $H_{i}(\operatorname{Cyl}_{f},N)=0$ for $i\neq q+1$, and $\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}([N,f]))$ is defined as a form on $\ker(H_{q}(f))\cong H_{q+1}(\operatorname{Cyl}_{f},N)$, which is equipped with the distinguished s-basis from Definition 2.50, see [Wall-SCM, Lemmas 2.3 and 5.1, Theorem 5.6]. Since $f$ is a homotopy equivalence, $H_{q+1}(\operatorname{Cyl}_{f},N)$ is trivial and $\tau(\operatorname{Cyl}_{f},N)=\tau(f)$ by Lemma 2.53. So by Lemma 2.51 the transition matrix between the canonical and the distinguished s-basis of $H_{q+1}(\operatorname{Cyl}_{f},N)$ has Whitehead torsion $(-1)^{q}\tau(f)$. In $L^{\mathrm{s}}_{n}(\mathbb{Z}G,w)$ the same element $\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}([N,f]))$ is also represented by a standard hyperbolic form with a basis such that the transition matrix between the standard and the chosen basis has Whitehead torsion $(-1)^{q}\tau(f)$.
@@
The image of $\pi(\tau(f))$ in $L^{\mathrm{s}}_{n}(\mathbb{Z}G,w)$ is represented by a standard hyperbolic form with any basis that has the following property: if $x$ denotes the Whitehead torsion of the transition matrix between the standard and the chosen basis, then $x\in\mathcal{J}_{n}(G,w)$ and $\pi(x)=\pi(\tau(f))$ (see [Shaneson-GxZ, p. 312]). In particular, the representative of $\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}([N,f]))$ constructed above also represents $\varrho(\widehat{\tau}([N,f]))$, noting that $\pi((-1)^{q}\tau(f))=\pi(\tau(f))$, because $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is $2$-torsion. Hence the square commutes if $n$ is even.
@@
Next consider the case when $n=2q+1$ is odd. Let $U$ denote a tubular neighbourhood of a disjoint union of embeddings $S^{q}\rightarrow N$ representing a generating set of $\ker(\pi_{q}(f))$ and let $N_{0}=N\setminus\operatorname{Int}U$. We identify $H_{q}(\partial U)$ with the standard hyperbolic form, then $H_{q+1}(U,\partial U)$ (more precisely, its image under the boundary map) corresponds to the standard lagrangian. We can assume that $f\big|_{U}$ is constant, so the restrictions of $f$ determine a map of pairs $(f\big|_{N_{0}},f\big|_{\partial U})\colon(N_{0},\partial U)\rightarrow(M,\ast)$. Then $\ker(H_{q+1}(N_{0},\partial U)\rightarrow H_{q+1}(M,\ast))\cong H_{q+2}(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}})$ determines another lagrangian in $H_{q}(\partial U)$. We equip $H_{q}(\partial U)$ and $H_{q+1}(U,\partial U)$ with their standard bases, and $H_{q+2}(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}})$ with the distinguished s-basis from Definition 2.50. Then $\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}([N,f]))$ is represented by the formation $(H_{q}(\partial U);H_{q+1}(U,\partial U),H_{q+2}(\operatorname{Cyl}_{f|_{N_{0}}},N_{0}\cup\operatorname{Cyl}_{f|_{\partial U}}))$, see [Wall-SCM, Section 6]. Since $f$ is a homotopy equivalence, we can take the empty generating set for $\ker(\pi_{q}(f))$. Then $H_{q}(\partial U)=0$, so we have the trivial formation, with the standard basis on the ambient form and on the first lagrangian, and the distinguished s-basis on the second lagrangian, $H_{q+2}(\operatorname{Cyl}_{f},N)$. By Lemma 2.51 applied with $C_{*}=C_{*}(\operatorname{Cyl}_{f},N)$, the transition matrix between the canonical and the distinguished s-basis has Whitehead torsion $(-1)^{q+1}\tau(\operatorname{Cyl}_{f},N)$, which by Lemma 2.53 equals $(-1)^{q+1}\tau(f)$.
@@
In $L^{\mathrm{s}}_{n}(\mathbb{Z}G,w)$ the same element $\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}([N,f]))$ is also represented by a formation on the standard hyperbolic form given by the standard lagrangians, such that the ambient form and the first lagrangian are equipped with their standard bases, and for the second lagrangian the transition matrix between the standard and the chosen basis has Whitehead torsion $(-1)^{q+1}\tau(f)$.
The image of $\pi(\tau(f))$ in $L^{\mathrm{s}}_{n}(\mathbb{Z}G,w)$ is represented by a formation on the standard hyperbolic form given by the standard lagrangians, such that the ambient form and the first lagrangian are equipped with their standard bases, and if $x$ denotes the Whitehead torsion of the transition matrix between the standard and the chosen basis of the second lagrangian, then $x\in\mathcal{J}_{n}(G,w)$ and $\pi(x)=\pi(\tau(f))$ (see [Shaneson-GxZ, p. 310]). Again we use that we can ignore signs because $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is $2$-torsion. So the representative of $\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}([N,f]))$ constructed above also represents $\varrho(\widehat{\tau}([N,f]))$, showing that the square also commutes if $n$ is odd. ∎
@@
Let $x\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$. If $x\in\operatorname{Im}\psi$, then there is a $\operatorname{CAT}$ $n$-manifold NN and a homotopy equivalence $f\colon N\to M$ such that $\pi(\tau(f))=x$.
@@
If $x=\psi(z)$ for some $z\in L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)$, then let $N,f]:=W_{\mathrm{h}}(z)$. By Proposition 3.6 we have $\pi(\tau(f))=\widehat{\tau}([N,f])=\widehat{\tau}(W_{\mathrm{h}}(z))=\psi(z)=x$. ∎
@@
$\operatorname{Im}\psi\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$.
@@
By the exactness of the sequence (3.2), and since $\operatorname{Id}_{M}$ has vanishing surgery obstruction, $\operatorname{Im}\psi=\ker\varrho\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$. ∎
Corollary 3.11.
$\operatorname{Im}\widehat{\tau}=\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$.
Proof.
The commutativity of the square in Proposition 3.7 implies that $\operatorname{Im}\widehat{\tau}\subseteq\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$. For the other direction assume that $x\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ and $\varrho(x)=\sigma_{\mathrm{s}}(y)$ for some $y\in\mathcal{N}(M)$. Then $\sigma_{\mathrm{h}}(y)=F\circ\sigma_{\mathrm{s}}(y)=F\circ\varrho(x)=0$, so $y=\eta_{\mathrm{h}}(v)$ for some $v\in\mathcal{S}^{\mathrm{h}}(M)$. Then we have $\varrho(x-\widehat{\tau}(v))=\varrho(x)-\sigma_{\mathrm{s}}(\eta_{\mathrm{h}}(v))=\varrho(x)-\sigma_{\mathrm{s}}(y)=0$, so $x-\widehat{\tau}(v)=\psi(z)$ for some $z\in L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)$. So by Proposition 3.6 $x=\widehat{\tau}(v)^{z}=\widehat{\tau}(v^{z})\in\operatorname{Im}\widehat{\tau}$. ∎
@@
There is a common generalisation of the two realisation techniques used in Section 3. Let $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$ denote the surgery obstruction groups defined in [Kreck-preprint, Section 4]. Similarly to $L_{n+1}^{\mathrm{s}}(\mathbb{Z}G,w)$, the elements of $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$ are represented by based nonsingular forms/formations, but, unlike the elements of $L_{n+1}^{\mathrm{s}}(\mathbb{Z}G,w)$, they are not assumed to be simple. Thus there is a natural forgetful map $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)\to L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)$ and $L_{n+1}^{\mathrm{s}}(\mathbb{Z}G,w)$ is the kernel of a map $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)\to\mathcal{J}_{n}(G,w)\leq\operatorname{Wh}(G,w)$ (see [Kreck-preprint]). There is also an action of $\operatorname{Wh}(G,w)$ on $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$ (which determines a map $\operatorname{Wh}(G,w)\to L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$ by acting on 0) given by changing the basis; this action is transitive on the fibres of the map $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)\to L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)$.
@@
Wall’s construction for realising elements of $L_{n+1}^{\mathrm{s}}(\mathbb{Z}G,w)$  [Wall-SCM] can also be applied to the group $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$, but it will only produce homotopy equivalences, not simple homotopy equivalences. Thus there is a map
$W_{\mathrm{s},\tau}\colon L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)\rightarrow\mathcal{S}^{\mathrm{h}}_{\mathrm{sCob}}(M),$
where $\mathcal{S}^{\mathrm{h}}_{\mathrm{sCob}}(M)$ is the set of homotopy equivalences $f\colon N\rightarrow M$ (as in $\mathcal{S}^{\mathrm{h}}(M)$), considered up to $s$-cobordism over $M$ (as in $\mathcal{S}^{\mathrm{s}}(M)$). The Whitehead torsion of the resulting homotopy equivalence is given by the previously mentioned map $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)\to\mathcal{J}_{n}(G,w)$.
The realisation part of the $s$-cobordism theorem determines a map
$W_{\operatorname{Wh}}\colon\operatorname{Wh}(G,w)\to\mathcal{S}^{\mathrm{hCob}}_{\mathrm{sCob}}(M),$
@@
where $\mathcal{S}^{\mathrm{hCob}}_{\mathrm{sCob}}(M)$ is the set of $h$-cobordisms $(W;M,N)$ considered up to $s$-cobordism rel. $M$. This is a special case of the above construction via the maps $\operatorname{Wh}(G,w)\to L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$ and $\mathcal{S}^{\mathrm{hCob}}_{\mathrm{sCob}}(M)\rightarrow\mathcal{S}^{\mathrm{h}}_{\mathrm{sCob}}(M)$. The Whitehead torsion of the homotopy equivalence induced by the $h$-cobordism we get is of course in $\mathcal{I}_{n+1}(G,w)$. The realisation of $L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)$, and the map $W_{\mathrm{h}}$, can also be defined using $W_{\mathrm{s},\tau}$, by first choosing a basis for the form/formation representing an element of $L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)$, equivalently, choosing a lift in $L_{n+1}^{\mathrm{s},\tau}(\mathbb{Z}G,w)$. As the choice of basis is not unique, we obtain a homotopy equivalence only up to $h$-cobordism (i.e. an element of $\mathcal{S}^{\mathrm{h}}(M)$), and its Whitehead torsion (the value of $\widehat{\tau}$) is only well-defined in the quotient $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$. See Remark 4.23 for a diagram summarising the different realisation maps, and their relationships to the simple homotopy manifold sets.